OpenStudy (anonymous):

MEDAL AND FAN!! A hyperbola with a horizontal transverse axis contains the point at (4, 3). The equations of the asymptotes are y - x = 1 and y + x = 1. Write an equation for the hyperbola.

OpenStudy (anonymous):

@ganeshie8 pls help

OpenStudy (dan815):

y=x+1, y=-x+1

OpenStudy (dan815):



OpenStudy (anonymous):

ok

OpenStudy (dan815):

pint 4,3 contained so

OpenStudy (dan815):